undeadknight26 (undeadknight26):

Let f(x) = 5x + 12. Find f-1(x).

undeadknight26 (undeadknight26):

So far i have: f(x) = 5x + 12 y = 5x + 12 x = 5y + 12 sqrtx = sqrt5y + 12

OpenStudy (anonymous):

Find y there in terms of x..

undeadknight26 (undeadknight26):

I have (x - 12)/(5)

OpenStudy (anonymous):

I mean from this equation : x = 5y + 12 Finding y in terms of x.. \(x - 12 = 5y\)

undeadknight26 (undeadknight26):

Or would it be (5)/x - 12? Im quite sure its (x-12)/(5)

OpenStudy (anonymous):

What else, replace y with \(f^{-1}(x)\) there..

OpenStudy (acxbox22):

you are right

undeadknight26 (undeadknight26):

I am?

undeadknight26 (undeadknight26):

its (x-12)/(5)? @acxbox22

OpenStudy (anonymous):

Yes, what you have is what you are looking for.. :)

OpenStudy (anonymous):

Yep..

OpenStudy (acxbox22):

yes that is the inverse y=(x-5)/12

undeadknight26 (undeadknight26):

Thank you very much!

OpenStudy (anonymous):

Yes, but why are you missing LHS?? Do write LHS there with your answer.. :)

undeadknight26 (undeadknight26):

LHS?

OpenStudy (anonymous):

You have: \[y = \frac{x-12}{5} \implies f^{-1}(x) = \frac{x-12}{5}\]

OpenStudy (anonymous):

LHS : Left Hand Side RHS : Right Hand Side.. An equation must contain LHS and RHS..

OpenStudy (acxbox22):

true..i forgot that

undeadknight26 (undeadknight26):

I seem to have not learned that yet...Will do because maybe i can get extra credit!

OpenStudy (anonymous):

I am just saying, that you should not (x-12)/5 alone.. Do mention what is it equal to.. It is actually : y = ..........